OpenStudy (anonymous):

a continuous compounding question with diff. equation and initial conditions. don't know where to start. please help.

OpenStudy (xapproachesinfinity):

State your problem properly as it was stated in your hw!

OpenStudy (anonymous):

You have $1000 with which to open an account which you plan also to add $1000 per year. All funds in the account will earn 10% annual interest compounded continuously. If the added deposits are also credited to your account continuously, the number of dollars x in your account at time t (yrs) will satisfy the initial value problem. Diff. equeation dx/dt = 1000 + 0.10x initial condition: x(0) = 1000 (a) solve the initial value problem for x as a function of t (b) about how many yrs will it take for the amount in your acount to reach 100,000

OpenStudy (aum):

\[ \frac{dx}{dt} = 1000 + 0.10x \\ \frac{dx}{1000+0.10x} = dt \\ \int \frac{dx}{1000+0.10x} = \int dt \\ \frac{\ln(1000+0.10x)}{0.10} = t \\ \ln(1000+0.10x) = 0.1t \\ (1000+0.10x) = e^{0.1t} \\ 0.1x = e^{0.1t} - 1000 \\ x = 10e^{0.1t} - 10,000 + C \\ t = 0, x = 1000 \\ 1000 = 10 - 10,000 + C \\ C = 10,990 \\ x = 10e^{0.1t} - 10,000 + 10,990 \\ x = 10e^{0.1t} + 990 \\ \]